OpenStudy (anonymous):

Solve 0=x^3-4x2-7x+28 and tell if it is extraneous

OpenStudy (anonymous):

@sepeario

OpenStudy (anonymous):

Factor by grouping \[0=(x^3-4x^2)+(-7x+28)\] Factor the GCF out of each ( ) so that you have the same expression left in both

OpenStudy (anonymous):

So, I got (x^2-7)(x-4)

OpenStudy (anonymous):

yes then set each of those equal to 0 and solve for x

OpenStudy (anonymous):

x=-7,7,4

OpenStudy (anonymous):

the 4 is right \[x^2-7=0\] \[x^2=7\] \[x=\pm \sqrt7\]

OpenStudy (anonymous):

Is it extraneous?

OpenStudy (sepeario):

Do you know the meaning of an extraneous solution?